# Applications of Spiking Neural Networks in Visual Place Recognition

### v3

@@ I Introduction
Visual place recognition is a critical component in robot localization tasks such as loop closure detection in Simultaneous Localization and Mapping (SLAM), and global re-localization of mobile robots [19, 13, 18, 16]. It is also relevant to image retrieval and landmark recognition tasks [15, 13]. Within robot navigation, VPR can minimize localization errors by recognizing previously visited places and updating the map of the environment despite appearance changes [15, 13, 20], which enables mobile robots to operate over extended periods.
@@
In the following sections, we will delve into the related works (Section II) to provide context for spiking neural networks and visual place recognition, discuss our methodology (Section III), and experimental setup (Section IV), present our results with analysis (Section V), and propose future research directions (Section VI), offering a thorough exploration of spiking neural networks for scalable visual place recognition.
@@ II Related works
II-A Neuromorphic Computing and Spiking Neural Networks
@@
II-B Spiking Neural Networks for Robot Localization
@@
II-C Visual Place Recognition
@@ III-A Preliminaries
Network structure: Each expert module, as illustrated in Figure 2, consists of a three-layer network architecture: 1) The input layer transforms each input image into Poisson-distributed spike trains via pixel-wise rate coding. The number of input neurons $K_{P}$ corresponds to the number of pixels in the input image: $K_{P}=W\times H$, where $W$ and $H$ correspond to the width and height of the input image respectively. 2) The $K_{P}$ input neurons are fully connected to $K_{E}$ excitatory neurons. Each excitatory neuron learns to represent a particular place, and a high firing rate of an excitatory neuron indicates high similarity between the learned and presented place. Note that multiple excitatory neurons can learn the same place. 3) Each excitatory neuron connects to exactly one of the $K_{I}$ inhibitory neurons. These inhibitory neurons inhibit all excitatory neurons except the excitatory neuron it receives a connection from. This enables lateral inhibition, resulting in a winner-takes-all system, that generates competition among excitatory neurons for effective learning.
Neuronal dynamics: The spiking neurons are modeled through their neuronal dynamics using the Leaky-Integrate-and-Fire (LIF) model [5]. The neuronal dynamics of excitatory neurons’ internal voltage is as follows:
@@
Network connections: The synaptic weights between input neurons and excitatory neurons, $W_{PE}$, are modeled via conductance changes to ensure stable network activity. When a synapse receives a presynaptic spike, the synaptic conductance is instantaneously increased by its synaptic weight and the synaptic conductance otherwise decays exponentially, as modeled by:
@@
Weight updates: The synaptic weights between the inhibitory and excitatory neurons are defined with constant synaptic weights, $W_{EI}$ and $W_{IE}$. The synaptic weights between input neurons and excitatory neurons, $W_{PE}$, are learned via the biologically inspired unsupervised learning mechanism Spike-Timing-Dependent-Plasticity (STDP). The weights are increased if the presynaptic spike occurs before a postsynaptic spike, and decreased otherwise. The synaptic weight change $\Delta w_{pe}$ after an input neuron $p$ receives a postsynaptic spike from an excitatory neuron $e$ is defined by:
@@
Local regularization of excitatory neurons: To prevent individual excitatory neurons from dominating the response, homeostasis is implemented through an adaptive neuronal threshold. The voltage threshold of the excitatory neurons is increased by a constant $\Theta$ after the neuron fires a spike, otherwise the voltage threshold decreases exponentially. We note that the homeostasis provides regularization only on the local, expert-specific scale, not on the global modular-level scale.
Neuronal assignment: The network training encourages the network to discern the different patterns (i.e. places) that were presented during training. As the training is unsupervised, one needs to assign each of the $K_{E}$ excitatory neurons to one of the $L$ training places ($K_{E}\gg L$). Following [73], we record the number of spikes $S_{e,l}$ of the $e$-th excitatory neuron when presented with an image of the $l$-th place. The highest average response of the neurons to place labels across the local training data is then used for the assignment $A_{e}$, such that neuron $K_{e}$ is assigned to place $l^{*}$ if:
@@ III-B Modular Scheme
Modular network structure: The previous section described how to train individual spiking networks following [73, 59]. In the following sections, we present our novel modular spiking network, which consists of a set of =$\mathcal{M}=\{M_{1},\dots,M_{i},\dots,M_{N}\}$ expert modules. The $i$-th module is tasked to learn the places contained in non-overlapping subsets $R_{i}\in\mathcal{R}$ of the reference database $\mathcal{R}$, whereby
@@
Modular place matching decision: At deployment time, the query image $q$ is provided as input to all modules in parallel. The predicted place of the Modular SNN, $\hat{l}_{M}$, is obtained by considering the spike outputs of all modules, rather than just a single module as in Eq. (5). We thus refine Eq. (5) to integrate the contributions of all $N$ modules:
@@ III-D Ensemble of Modular SNNs
Ensemble network structure: In this section, we introduce ensembles of Modular SNNs. The purpose of these ensembles is to improve robustness and generalization ability. The main idea is that each place $l$ is represented by multiple complementary ensemble members. Specifically, the ensemble is represented as a set of =$\mathcal{E}=\{E_{1},\dots,E_{m},\dots,E_{M}\}$ homogeneous ensemble members (i.e. their network architecture is the same), where each member is an independent Modular SNN. The ensemble members are all trained in parallel on the entire reference database $\mathcal{R}$ (Figure 1).
@@
Creation of distance matrix: To compute the single-frame distance matrix for the Ensemble of Modular SNNs, $D_{\text{single}}$, we compute the cumulative spike activity for all neurons assigned to each place label for each query image $q$. This contrasts with taking only the maximum that is considered as the prediction in Eq. 10. Specifically, given a query image $q$, the $q$-th column of the single-frame similarity matrix $S_{\text{single}}$ is defined as:
@@ IV-A Implementation Details
We implement our SNNs using the Brian2 simulator. We publicly release our code at: https://github.com/QVPR/VPRSNN. Each SNN module contains $K_{E}=400$ excitatory neurons, $K_{I}=400$ inhibitory neurons, and $K_{P}=W\times H$ input neurons, where the input image height and width is $28$ pixels in each dimension. We select the network’s hyperparameters as follows: For the number of training epochs, we use a fixed value of $30$ epochs for all datasets. For the threshold value $\theta$ (Eq. (8)) to detect the hyperactive neurons for each individual Modular SNN evaluated on a dataset, we draw the threshold value from a uniformly random distribution with $U(40,100)$ for small-scale datasets ($<1000$ places) and $U(600,800)$ for large-scale datasets ($>=1000$ places). The threshold value of each Modular SNN within an Ensemble of Modular SNNs evaluated on a dataset, is chosen separately based on this uniformly random selection process. For our Ensemble of Modular SNNs in the results section, Section V, we used three or five ensemble members, with the exact number specified in each part.
@@ IV-B Datasets
We evaluate our work using several datasets that cover seasonal changes [54, 57], differences in the time of day [55, 56, 57], as well as rural [56, 54], suburban [58] and urban [55, 57] environments with additional challenges due to occlusions and glare [58]. We now briefly describe these datasets, and provide sample images in Figure 3. Our code base [^2] https://github.com/QVPR/VPRSNN provides details on how these datasets are used.
@@
The Nordland dataset [54]: captures a 728 km train path in Norway recorded in spring, summer, fall and winter. As commonly done in the literature [128, 129, 130], the data segments where the speed of the train is below 15km/h is removed using the provided GPS data. We used the Nordland dataset configured into two different sets: 1) Reference: Fall; query: Summer (also referred to as Nordland FS); 2) Reference: Spring; query: Winter (also referred to as Nordland SW). We considered every image in a traverse as a distinct place, obtaining 27575 places for each traverse.
@@
The Oxford RobotCar dataset [55]: has over 100 traverses captured in Oxford city and it is recorded at varying conditions including different time of the day, and different seasons. As done in prior works [131, 59], we selected the front left stereo frames from the Rain (2015-10-29-12-18-17) traverse as the reference dataset and from the Dusk (2014-11-21-16-07-03) traverse as the query dataset. For each traverse, we sampled approximately one image per meter, resulting in 3800 places.
@@
The SFU-Mountain dataset [56]: has more than 8 hours of trail driving in Burnaby Mountain, British Columbia Canada, using the Clearpath Husky robot covering sunny, rainy and snowy conditions [56]. Following prior works [132, 133], we used the front right stereo frames from the dry traverse for reference and from the dusk traverse for query. We considered each image in the traverse as a different place, and used the author’s place sampling configurations [56], using the entire dataset where each traverse contains 375 places.
@@
The Synthia Night-to-Fall dataset [57]: is a synthetic dataset that was initially designed for semantic scene understanding in a city-like driving scenario. In our approach, following the methodology described in [134], we used the SEQS-04 foggy (reference) and nighttime (query) traverses. Segments in which the vehicle remained stationary were excluded, and we sampled approximately one frame per meter resulting in 250 places for each traverse.
@@
The St Lucia dataset [58]: comprises several traverses along a route within the St Lucia suburb of Brisbane. For our experiments, we employed a traverse conducted during the early morning (190809-0845) as the reference traverse and another in the afternoon (180809-1545) as the query traverse. We omitted the segments where the vehicle was at rest and sampled places approximately every 15 meters. We used only the unique places from the reference traverse, obtaining 500 places. For the query traverse, we included sections where places were visited multiple times within the same traverse, obtaining 1037 places.
@@ IV-E Baseline Methods
Sum-of-Absolute-Differences (SAD) [41]: is a simple baseline technique which computes the pixel-wise difference between each query image and all reference images. While the sum-of-absolute differences method is simple, this method is not very robust to drastic varying conditions such as changes in lighting, viewpoint, and seasonal changes.
@@
NetVLAD [49]: aggregates local image descriptors with learnable aggregation weights by employing the VLAD technique [102] to create a fixed-size global descriptor. This model has a VGG16 [135] backbone pretrained on ImageNet [136]. The NetVLAD layer, which is the pooling layer of the network, is trained on the Google Landmarks [137], Mapillary Street Level Sequences [138], and Pittsburgh [139] datasets separately. As NetVLAD is trained on urban environments, the model might not generalize well to completely different types of environments such as rural areas.
@@
DenseVLAD [48]: utilizes densely sampled SIFT [140] image descriptors, and aggregates these features using VLAD [102]. While DenseVLAD is very robust to high illumination, some limitations of DenseVLAD include lack of robustness to occlusions, very dark conditions with limited dynamic range, and rural areas with vegetation.
@@
AP-GeM [50]: employs the Generalized-Mean pooling layer (GeM) [141] and uses a listwise loss formulation that directly optimizes for the Average Precision (AP) performance metric. This model uses a CNN-based backbone pretrained on ImageNet [136] to extract feature representations and aggregates it into a compact representation. We used three variations of AP-GeM; a ResNet50 backbone [142] trained on Landmarks-clean [143] dataset, a ResNet101 backbone [142] trained on Landmarks-clean [143] dataset, and a Res101 backbone [142] trained on Google-Landmarks [137] Dataset. Similar to NetVLAD, AP-GeM generalization is reliant on the type of environment used for training.
@@
GCL [51]: is trained via a Generalized Contrastive Loss (GCL) using graded similarity labels for image pairs. We trained the last two layers of the network on the Mapillary Street-Level Sequences (MSLS) dataset [138] using a VGG16 backbone [135] pretrained on ImageNet [136] with GeM [141] as the pooling layer.
@@
CosPlace [52]: uses a classification framework to train the model. This approach splits the training dataset into square geographical cells using the ground truth data. During training, the network iterates over CosPlace Groups, which are non-overlapping classes that are grouped together. The network uses the Large Margin Cosine Loss (LCML) [144] with a fully connected layer, that is only present at training, for each group of the dataset. The network consists of a CNN backbone (a VGG-16 backbone), a GeM pooling layer, and a fully connected output layer. At inference, the model outputs compact discriminative feature descriptors.
@@
MixVPR [53]: is a global feature aggregation method, that takes the feature maps of a intermediate layers of a CNN-based backbone (a ResNet backbone), that are processed by a cascade of Feature Mixer layers, comprised of multi-layer perceptrons, that provide each element of the feature map with global relationships to all other elements. MixVPR is a SOTA VPR technique across multiple VPR benchmark datasets.

### v4

@@ I Introduction
VPR is a critical component in robot localization tasks such as loop closure detection in Simultaneous Localization and Mapping (SLAM), and global re-localization of mobile robots [19, 13, 18, 16]. It is also relevant to image retrieval and landmark recognition tasks [15, 13]. Within robot navigation, VPR can minimize localization errors by recognizing previously visited places and updating the map of the environment despite appearance changes [15, 13, 20], which enables mobile robots to operate over extended periods.
@@
The rest of this article is organized as follows. In Section II, we will delve into the related works to provide context for SNNs and VPR; in Section III, we will discuss our methodology; experimental setup is given in Section IV; Section V presents our results with analysis; and finally, Section VI concludes this article.
@@ II Related works
II-A Neuromorphic Computing and SNNs
@@
II-B SNNs for Robot Localization
@@
II-C Visual Place Recognition (VPR)
@@ III-A Preliminaries
III-A1 Network Structure
Each expert module, as illustrated in Figure 2, consists of a three-layer network architecture:
  1. i.
    The input layer transforms each input image into Poisson-distributed spike trains via pixel-wise rate coding. The number of input neurons $K_{P}$ corresponds to the number of pixels in the input image: $K_{P}=W\times H$, where $W$ and $H$ correspond to the width and height of the input image respectively.
  2. ii.
    The $K_{P}$ input neurons are fully connected to $K_{E}$ excitatory neurons. Each excitatory neuron learns to represent a particular place, and a high firing rate of an excitatory neuron indicates high similarity between the learned and presented place. Note that multiple excitatory neurons can learn the same place.
  3. iii.
    Each excitatory neuron connects to exactly one of the $K_{I}$ inhibitory neurons. These inhibitory neurons inhibit all excitatory neurons except the excitatory neuron it receives a connection from. This enables lateral inhibition, resulting in a winner-takes-all system, that generates competition among excitatory neurons for effective learning.
III-A2 Neuronal Dynamics
The spiking neurons are modeled through their neuronal dynamics using the Leaky-Integrate-and-Fire (LIF) model [5]. The neuronal dynamics of excitatory neurons’ internal voltage is as follows:
@@ III-A2 Neuronal Dynamics
III-A3 Network Connections
The synaptic weights between input neurons and excitatory neurons, $W_{PE}$, are modeled via conductance changes to ensure stable network activity. When a synapse receives a presynaptic spike, the synaptic conductance is instantaneously increased by its synaptic weight and the synaptic conductance otherwise decays exponentially, as modeled by:
@@ III-A3 Network Connections
III-A4 Weight Updates
The synaptic weights between the inhibitory and excitatory neurons are defined with constant synaptic weights, $W_{EI}$ and $W_{IE}$. The synaptic weights between input neurons and excitatory neurons, $W_{PE}$, are learned via the biologically inspired unsupervised learning mechanism Spike-Timing-Dependent-Plasticity (STDP). The weights are increased if the presynaptic spike occurs before a postsynaptic spike, and decreased otherwise. The synaptic weight change $\Delta w_{pe}$ after an input neuron $p$ receives a postsynaptic spike from an excitatory neuron $e$ is defined by:
@@ III-A4 Weight Updates
III-A5 Local Regularization of Excitatory Neurons
To prevent individual excitatory neurons from dominating the response, homeostasis is implemented through an adaptive neuronal threshold. The voltage threshold of the excitatory neurons is increased by a constant $\Theta$ after the neuron fires a spike, otherwise the voltage threshold decreases exponentially. We note that the homeostasis provides regularization only on the local, expert-specific scale, not on the global modular-level scale.
III-A6 Neuronal Assignment
The network training encourages the network to discern the different patterns (i.e. places) that were presented during training. As the training is unsupervised, one needs to assign each of the $K_{E}$ excitatory neurons to one of the $L$ training places ($K_{E}\gg L$). Following [73], we record the number of spikes $S_{e,l}$ of the $e$-th excitatory neuron when presented with an image of the $l$-th place. The highest average response of the neurons to place labels across the local training data is then used for the assignment $A_{e}$, such that neuron $K_{e}$ is assigned to place $l^{*}$ if:
@@ III-B Modular Scheme
III-B1 Modular Network Structure
The previous section described how to train individual spiking networks following [73, 59]. In the following sections, we present our novel modular spiking network, which consists of a set of =$\mathcal{M}=\{M_{1},\dots,M_{i},\dots,M_{N}\}$ expert modules. The $i$-th module is tasked to learn the places contained in non-overlapping subsets $R_{i}\in\mathcal{R}$ of the reference database $\mathcal{R}$, whereby
@@ III-B1 Modular Network Structure
III-B2 Modular Place Matching Decision
At deployment time, the query image $q$ is provided as input to all modules in parallel. The predicted place of the Modular SNN, $\hat{l}_{M}$, is obtained by considering the spike outputs of all modules, rather than just a single module as in Eq. (5). We thus refine Eq. (5) to integrate the contributions of all $N$ modules:
@@ III-D Ensemble of Modular SNNs
III-D1 Ensemble Network Structure
In this section, we introduce ensembles of Modular SNNs. The purpose of these ensembles is to improve robustness and generalization ability. The main idea is that each place $l$ is represented by multiple complementary ensemble members. Specifically, the ensemble is represented as a set of =$\mathcal{E}=\{E_{1},\dots,E_{m},\dots,E_{M}\}$ homogeneous ensemble members (i.e. their network architecture is the same), where each member is an independent Modular SNN. The ensemble members are all trained in parallel on the entire reference database $\mathcal{R}$ (Figure 1).
@@ III-D2 Ensemble Place Matching Decision
III-D3 Creation of Distance Matrix
To compute the single-frame distance matrix for the Ensemble of Modular SNNs, $D_{\text{single}}$, we compute the cumulative spike activity for all neurons assigned to each place label for each query image $q$. This contrasts with taking only the maximum that is considered as the prediction in Eq. 10. Specifically, given a query image $q$, the $q$-th column of the single-frame similarity matrix $S_{\text{single}}$ is defined as:
@@ IV-A Implementation Details
We implement our SNNs using the Brian2 simulator. We publicly release our code online222[Online]. Available: https://github.com/QVPR/VPRSNN. Each SNN module contains $K_{E}=400$ excitatory neurons, $K_{I}=400$ inhibitory neurons, and $K_{P}=W\times H$ input neurons, where the input image height and width is $28$ pixels in each dimension. We select the network’s hyperparameters as follows: For the number of training epochs, we use a fixed value of $30$ epochs for all datasets. For the threshold value $\theta$ (Eq. (8)) to detect the hyperactive neurons for each individual Modular SNN evaluated on a dataset, we draw the threshold value from a uniformly random distribution with $U(40,100)$ for small-scale datasets ($<1000$ places) and $U(600,800)$ for large-scale datasets ($>=1000$ places). The threshold value of each Modular SNN within an Ensemble of Modular SNNs evaluated on a dataset, is chosen separately based on this uniformly random selection process. For our Ensemble of Modular SNNs in the results section, Section V, we used three or five ensemble members, with the exact number specified in each part.
@@ IV-B Datasets
We evaluate our work using several datasets that cover seasonal changes [54, 57], differences in the time of day [55, 56, 57], as well as rural [56, 54], suburban [58] and urban [55, 57] environments with additional challenges due to occlusions and glare [58]. We now briefly describe these datasets, and provide sample images in Figure 3. Our code base333[Online]. Available: https://github.com/QVPR/VPRSNN provides details on how these datasets are used.
IV-B1 The Nordland dataset [54]
@@
captures a 728 km train path in Norway recorded in spring, summer, fall and winter. As commonly done in the literature [128, 129, 130], the data segments where the speed of the train is below 15km/h is removed using the provided GPS data. We used the Nordland dataset configured into two different sets:
  1. i.
    Reference: Fall; query: Summer (also referred to as Nordland FS);
  2. ii.
    Reference: Spring; query: Winter (also referred to as Nordland SW).
We considered every image in a traverse as a distinct place, obtaining 27575 places for each traverse.
IV-B2 The Oxford RobotCar dataset [55]
@@
has over 100 traverses captured in Oxford city and it is recorded at varying conditions including different time of the day, and different seasons. As done in prior works [131, 59], we selected the front left stereo frames from the Rain (2015-10-29-12-18-17) traverse as the reference dataset and from the Dusk (2014-11-21-16-07-03) traverse as the query dataset. For each traverse, we sampled approximately one image per meter, resulting in 3800 places.
IV-B3 The SFU-Mountain dataset [56]
@@
has more than 8 hours of trail driving in Burnaby Mountain, British Columbia Canada, using the Clearpath Husky robot covering sunny, rainy and snowy conditions [56]. Following prior works [132, 133], we used the front right stereo frames from the dry traverse for reference and from the dusk traverse for query. We considered each image in the traverse as a different place, and used the author’s place sampling configurations [56], using the entire dataset where each traverse contains 375 places.
IV-B4 The Synthia Night-to-Fall dataset [57]
@@
is a synthetic dataset that was initially designed for semantic scene understanding in a city-like driving scenario. In our approach, following the methodology described in [134], we used the SEQS-04 foggy (reference) and nighttime (query) traverses. Segments in which the vehicle remained stationary were excluded, and we sampled approximately one frame per meter resulting in 250 places for each traverse.
IV-B5 The St Lucia dataset [58]
@@
comprises several traverses along a route within the St Lucia suburb of Brisbane. For our experiments, we employed a traverse conducted during the early morning (190809-0845) as the reference traverse and another in the afternoon (180809-1545) as the query traverse. We omitted the segments where the vehicle was at rest and sampled places approximately every 15 meters. We used only the unique places from the reference traverse, obtaining 500 places. For the query traverse, we included sections where places were visited multiple times within the same traverse, obtaining 1037 places.
@@ IV-E1 Sum-of-Absolute-Differences (SAD) [41]
is a simple baseline technique which computes the pixel-wise difference between each query image and all reference images. While SAD method is simple, this method is not very robust to drastic varying conditions such as changes in lighting, viewpoint, and seasonal changes.
IV-E2 NetVLAD [49]
@@
aggregates local image descriptors with learnable aggregation weights by employing the VLAD technique [102] to create a fixed-size global descriptor. This model has a VGG16 [135] backbone pretrained on ImageNet [136]. The NetVLAD layer, which is the pooling layer of the network, is trained on the Google Landmarks [137], Mapillary Street Level Sequences [138], and Pittsburgh [139] datasets separately. As NetVLAD is trained on urban environments, the model might not generalize well to completely different types of environments such as rural areas.
IV-E3 DenseVLAD [48]
@@
utilizes densely sampled Scale Invariant Feature Transform (SIFT) [140] image descriptors, and aggregates these features using VLAD [102]. While DenseVLAD is very robust to high illumination, some limitations of DenseVLAD include lack of robustness to occlusions, very dark conditions with limited dynamic range, and rural areas with vegetation.
IV-E4 AP-GeM [50]
@@
employs the Generalized-Mean pooling layer (GeM) [141] and uses a listwise loss formulation that directly optimizes for the Average Precision (AP) performance metric. This model uses a CNN-based backbone pretrained on ImageNet [136] to extract feature representations and aggregates it into a compact representation. We used three variations of AP-GeM; a residual networks (ResNet)50 backbone [142] trained on Landmarks-clean [143] dataset, a ResNet101 backbone [142] trained on Landmarks-clean [143] dataset, and a ResNet101 backbone [142] trained on Google-Landmarks [137] Dataset. Similar to NetVLAD, AP-GeM generalization is reliant on the type of environment used for training.
IV-E5 Generalized Contrastive Loss (GCL) [51]
@@
is trained via a GCL using graded similarity labels for image pairs. We trained the last two layers of the network on the Mapillary Street-Level Sequences (MSLS) dataset [138] using a VGG16 backbone [135] pretrained on ImageNet [136] with GeM [141] as the pooling layer.
IV-E6 CosPlace [52]
@@
uses a classification framework to train the model. This approach splits the training dataset into square geographical cells using the ground truth data. During training, the network iterates over CosPlace Groups, which are non-overlapping classes that are grouped together. The network uses the Large Margin Cosine Loss (LCML) [144] with a fully connected layer, that is only present at training, for each group of the dataset. The network consists of a CNN backbone (a VGG-16 backbone), a GeM pooling layer, and a fully connected output layer. At inference, the model outputs compact discriminative feature descriptors.
IV-E7 MixVPR [53]
@@
is a global feature aggregation method, that takes the feature maps of a intermediate layers of a CNN-based backbone (a ResNet backbone), that are processed by a cascade of Feature Mixer layers, comprised of multi-layer perceptrons, that provide each element of the feature map with global relationships to all other elements. MixVPR is a state-of-the-art VPR technique across multiple VPR benchmark datasets.
IV-E8 Non-modular SNN [59]